OpenStudy (kittiwitti1):

A bacteria culture initially contains 250 cells and grows at a rate proportional to its size. After 2 hours, the population has increased to 600 cells. A) Find a function for the number of bacteria after t hours (FINISHED). B) Find the number of bacteria after 8 hours (FINISHED). C) Find the rate of growth after 8 hours. D) When will the population reach 100,000? My questions are: Does (C) involve derivatives? And is (D) related to (C), or can I use the normal exponential growth/decay model?

OpenStudy (jadeishere):

So (c) is asking you to find the rate of growth, I guess you knew that, but.. So, the formula for exponential growth \[y = a(1 + r)^{x}\] where a = initial amount r = rate of growth x = number of time intervals that have gone by

OpenStudy (kittiwitti1):

wot

OpenStudy (kittiwitti1):

So then, this scenario would be \(y=250(1+r)^{8}\)?

OpenStudy (jadeishere):

Yep :)

OpenStudy (jadeishere):

Do you have a graphing calculator?

OpenStudy (kittiwitti1):

Er, I'm not sure the professor wants me to use one. But I do have a graphing calculator

OpenStudy (jadeishere):

Answer to your first question; No it does not involve derivatives.

OpenStudy (kittiwitti1):

If I solved for \(r\) would that work as well?

OpenStudy (kittiwitti1):

o: Eh... but the entire chapter section talks about derivatives. :(

OpenStudy (jadeishere):

Do you know what a derivative is?

OpenStudy (jadeishere):

And yes, if you solved for 'r' it should work, if done correctly

OpenStudy (kittiwitti1):

http://prntscr.com/cxyijf This is the first page of the section.

OpenStudy (kittiwitti1):

Immediately started off with (implicit?) differentiation

OpenStudy (kittiwitti1):

Oh, oops, I meant "differentiation," not "derivatives" But yeah -

OpenStudy (jadeishere):

Um.. I'm not sure how to address differentiation

OpenStudy (kittiwitti1):

Yeah... I don't know either. But thanks for trying.

OpenStudy (holsteremission):

The first sentence can be translated into a differential equation. If the size of the culture at time \(t\) is \(y(t)\), then the growth rate is \(\dfrac{\mathrm dy}{\mathrm dt}=y'(t)\). You're told that the growth rate is proportional to the culture size, so there is some constant \(k\) for which \[y'(t)=ky(t)\]This is a separable ODE, so you can write \[\frac{\mathrm dy}{y}=k\,\mathrm dt\]then integrate and solve for the arbitrary constant of integration using the fact that the culture starts with \(250\) cells, so \(y(0)=250\). Meanwhile, you're also told that at \(t=2\), the culture has \(600\) cells, so \(y(2)=600\), and this is enough info to solve for \(k\). I assume this is what you did for parts (a) and (b). Part (c) is asking you to find the (instantaneous) growth rate at \(t=8\), which is a matter of computing \[y'(8)=ky(8)=\cdots\]

OpenStudy (kittiwitti1):

Thank you @HolsterEmission :) I actually found an explanation online at the same time as your reply haha